Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-12-15
ISS On-Orbit Status 12/15/07

Saturday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Tani except for housekeeping and voluntary work.

After wakeup and before breakfast, FE-2 Dan Tani again accessed the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and completing questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for later downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Dan wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

The crew conducted the regular weekly three-hour task of thorough station cleaning. ("Uborka", normally done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

In the Lab, FE-2 Dan Tani wrapped up his support of the CSLM-2 (Coarsening in Solid-Liquid Mixtures 2) experiment in the MSG (Microgravity Science Glovebox) facility, terminating SPU-3 (Sample Processing Unit #3) processing, transferring its data to the MLC (MSG Laptop Computer) and verifying them, then removing SPU-3, installing a new SPU (#11), and finally turning the payload off. Later, the MSG was also powered off. (CSLM-2 examines the kinetics of competitive particle growth within a liquid matrix. During this process, small particles shrink by losing atoms to larger particles, causing the larger particles to grow (coarsen) within a liquid lead/tin matrix. This study defined the mechanisms and rates of coarsening that govern the manufacture with metals from turbine blades to dental amalgam fillings. By all indications, the CSLM runs were nominal.)

In Node-2, with the BCAT-3 (Binary Colloidal Alloy Test-3) science payload running by itself since 12/13, the FE-2 performed his daily status check on the A31p laptop controlling the EarthKAM digital still camera, verifying on the last image taken that image focus and alignment remain in check. (The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour for the next few days. The EarthKAM camera will be required next week for post-EVA EMU glove photography.)

FE-1 Yuri Malenchenko performed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

At ~6:35am EST, the FE-1 had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop).

At ~9:15am, the crewmembers held their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

The crewmembers performed their regular 2.5-hr physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

A new task added to the voluntary 'job jar' list for CDR and FE-2 calls for finishing up the audit and consolidation of photo/TV items in their CTBs (Cargo Transfer Bags), started by Dan Tani yesterday.

Also in the discretionary 'job jar' for Dan for today were departure preparations for his end-of-increment cleanup. (Due to the extension of his station tenure to NET mid-January, the FE-2 will have more leisure in completing this clean-up.)

Working from his own 'time permitting' job list, Yuri was to perform the periodic collection & deletion of readings on the MOSFET (metal oxide semiconductor field-effect transistor) radiation sensor reader/display of the RBO-3-2 Matryoshka-R antroph-amorphous (human torso) "phantoms" located inside the station for sophisticated radiation studies, collecting radiation measurements every 15 minutes around the clock.

A second job item on Malenchenko's discretionary list for today was another KPT-3 session to make observations and take aerial KPT-3 photography of environmental conditions for Russia's Environmental Safety Agency (ECON) using the Nikon D1X digital camera with SIGMA 300-800mm telephoto lens. (KPT-3 photography has been a frequent earth observing experiment for ECON.)

MT Translation: Yesterday's planned MT (Mobile Transporter) translation from WS7 (Worksite 7) to WS4, erroneously reported here as completed, was deferred to no earlier than 12/20 (Thursday) due to a possible obstruction of MLI (Multi-Layered Insulation) on the NTA (Nitrogen Tank Assembly). The translation was to provide added protection of the TUS (Trailing Umbilical System) from MMOD (Micrometeoroid/Orbital Debris) between now and Flight 1E. Ground controllers are assessing imagery of the translation path and developing a forward plan.

BGA 1A Update: More testing was performed last night on the failed 1A BGA (Beta Gimbal Assembly). During a power switch sequence the BCDU 3A1 CP RBI (Battery Charge/Discharge Unit/CP Remote Bus Isolator), a fuse-type switch, did not trip open, but the RPCs (Remote Power Controllers) of both primary and redundant power feeds to the 1A ECU (Electronics Control Unit) tripped several times. At the conclusion of the test, the RPCs remained closed for about 3 hours, then tripped again. ECU 1A is now unpowered and BGA 1A is locked, with the motor off, i.e. in stable configuration. Engineering analysis of the dumped data continues, but no more troubleshooting is expected before the EVA-13.

Weekly Science Update (Expedition Sixteen -- Week 8)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue nominally in the PIRS module. The next memory card replacement activity is currently scheduled on 12/10.

ANITA: Completed.

BCAT (Binary Colloidal Alloy Test): The Binary Colloidal Alloy Test-3 experiment can now proceed, thanks to Astronauts Dr. Peggy Whitson and Dan Tani. Sample 3, the most critical of the samples (and the one closest to the critical point) was the first to be restarted to allow for the publication of the science data by Harvard University (Dr. David Weitz, PI, and Peter Lu, Doctorate Candidate). Operations (EarthKAM imaging) will continue for 4 days when parts of the BCAT-3 setup is needed for EVA activities. After the EVA BCAT-3 activities will continue for a total of 14 days.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan: Due to the slip of 1E, we will need to repeat your 2nd session at the appropriate time frame (within R -21 to R -14) to meet the experiment's requirements. This activity will be hard scheduled. Again, thank you for your participation.'

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): CSLM-2 SPU-3 34-hr soak is in progress and should be complete by 12/15/07.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

EPO (Educational Payload Operations): EPO is very satisfied with the EPO Demos that the crew conducted within the past week including the Sanitation on Station, ISS Tour / Living Space, and Newton's Laws Demo. The videos will be used to enhance existing education resources. 'We would like to thank the crew for their efforts in inspiring the next generation of explorers in science, technology, engineering, and mathematics.'

ETD (Eye Tracking Device): Second session for Yuri Malenchenko was performed nominally on 12/12.

Integrated Immune: In progress.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples will be downloaded on STS-122 (1E). Peggy performed successfully the Water Reservoirs Exchange activity on 12/12. That allows for additional dry-out of the EMCS (European Modular Cultivation System) Facility as preventive measure to the now delayed Clean-Up activities planned during 1E Stage..

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy: The Sleep PI appreciates you going 'above and beyond' with your extra sleep logging. Remaining for the Sleep experiment, you have monthly downloads, doffing the Actiwatch, and two more weeks of required sleep logging.' 'Dan: The Sleep team would appreciate if you could continue to sleep log so that the Sleep experiment's logging requirements (3 weeks during the mission) are met.'

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress. To be conducted tomorrow (12/2).

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

CEO (Crew Earth Observation): On-going.

CEO photo targets uplinked for today were Afar Rift Zone, Ethiopia (looking to the right of track as ISS approached the "elbow" formed by the Red Sea and Gulf of Aden - the area to the west of the elbow is the Afar Rift Zone (also known as the Afar Triangle). This geologically active region is a junction between the Arabian, Indian, and African tectonic plates - all of which are moving away from each other, producing frequent earthquakes and historically active volcanoes (such as Erta Ale in Ethiopia). Overlapping frames of the Afar Triangle area will provide a useful record of the rift area, particularly surface expressions of faults and lava flows), and Madrean Sky Islands (this target is located in the northern reaches of Mexico's Sierra Madre Occidental which boast some of the richest biodiversity anywhere in North America. It is a comprised of cool, moist, higher-altitude pine-oak forested mountain ranges that dot the hot, lower Sonoran and Chihuahuan deserts of southern Arizona and New Mexico and northwestern Mexico. These climatological islands are situated in remote and rugged areas and are vestiges of cooler, wetter periods during the ice ages. Nadir- or near-nadir imagery of the forested mountaintops was requested).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use